# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ A.3 Gronwall’s Inequality
Since $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ is a saddle point of
@@
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{w}_{z}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})-\mathbf{y}\right\|^{2},$ (56)
it follows that
$\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\end{bmatrix}=\begin{bmatrix}\nabla\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})^{\top}\overline{\mathbf{y}}\\ \nabla\mathcal{H}_{z}(\mathbf{X};\overline{\mathbf{w}}_{z})^{\top}\overline{\mathbf{y}},\end{bmatrix},$
where $\overline{\mathbf{y}}=\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})$. Also, we know that if $\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma,$ then
$\left\langle\mathbf{w}_{n}-\overline{\mathbf{w}}_{n},\nabla_{\mathbf{w}_{n}}\mathcal{L}(\mathbf{w}_{n},\mathbf{0})\right\rangle\geq 0.$
In this section, $\mathcal{J}_{z}\left(\mathbf{X};\mathbf{w}_{z}\right)$ and $\mathcal{J}_{n}\left(\mathbf{X};\mathbf{w}_{n}\right)$ denotes the Jacobian of $\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})$ and $\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})$, respectively. Then, the above inequality can be written as
$\left\langle\mathbf{w}_{n}-\overline{\mathbf{w}}_{n},\mathcal{J}_{n}\left(\mathbf{X};\mathbf{w}_{n}\right)^{\top}(\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})-\mathbf{y})\right\rangle\geq 0.$ (57)
Let $\mathbf{p}_{z}$ be a fixed unit norm vector and $\mathbf{u}(t)$ be the solution of
$\frac{d\mathbf{u}}{dt}=\nabla\mathcal{N}_{\overline{\mathbf{y}},\mathcal{H}_{z}}(\mathbf{u})=\mathcal{J}_{z}\left(\mathbf{X};\mathbf{u}\right)^{\top}\overline{\mathbf{y}},\mathbf{u}(0)=\mathbf{p}_{z}.$
Similar to the proof of Theorem 5.1, we next define $\eta$, $T_{\epsilon}$ and $B_{\epsilon}$. From Lemma 5.2, we know $\mathbf{u}(t)$ either converges to $\mathbf{0}$, or $\mathbf{u}(t)$ converges in direction to a non-negative KKT point of constrained NCF defined with respect to $\overline{\mathbf{y}}$ and $\mathcal{H}_{z}$. If $\mathbf{u}(t)$ converges to $\mathbf{0}$, then we define $\eta=\min(2,2\gamma).$ Then, for any $\epsilon\in(0,\eta/2)$, we define $B_{\epsilon}=\epsilon$ and choose $T_{\epsilon}$ large enough such that
$\|\mathbf{u}(T_{\epsilon})\|_{2}\leq B_{\epsilon}=\epsilon.$ (58)
Else, if $\mathbf{u}(t)$ does not converge to $\mathbf{0}$, then from Lemma B.6, there exists $\tilde{\eta}>0$ and $T$ such that $\|\mathbf{u}(t)\|_{2}\geq\tilde{\eta}$, for all $t\geq T$. We next define $\eta=\min(\tilde{\eta},1,2\gamma)$. Then, from Lemma B.6, for any $\epsilon\in(0,\eta/2)$ there exists $T_{\epsilon}\geq T$ and $B_{\epsilon}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(T_{\epsilon})\|_{2}\leq B_{\epsilon},\forall t\leq T_{\epsilon},$ (59)
where $\mathbf{u}^{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\overline{\mathbf{y}},\mathcal{H}_{z}}(\mathbf{u}).$
Additionally, since $T_{\epsilon}\geq T$, we have
$\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\tilde{\eta}\geq\eta.$ (60)
We define
@@
$\beta=\max\left(\sup\{\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})\|_{2}:\|\mathbf{w}_{n}\|_{2}=1\},\sup\{\|\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})\|_{2}:\|\mathbf{w}_{z}\|_{2}=1\}\right).$ (61)
and
$\alpha=\max\left(\sup\{\|\mathcal{J}_{n}(\mathbf{X};\mathbf{w}_{n})\|_{2}:\|\mathbf{w}_{n}\|_{2}=1\},\sup\{\|\mathcal{J}_{z}(\mathbf{X};\mathbf{w}_{z})\|_{2}:\|\mathbf{w}_{z}\|_{2}=1\}\right).$ (62)
Next, we define other $\epsilon$ dependent parameters which are useful in our proof. Let $\tilde{B}_{\epsilon}=B_{\epsilon}+\epsilon,\tilde{W}_{\epsilon}=\|\overline{\mathbf{w}}_{n}\|_{2}+\epsilon$, and $C_{\epsilon}=(1+2\epsilon\alpha\tilde{W}_{\epsilon}^{L-1}\beta\tilde{B}_{\epsilon}^{L}T_{\epsilon})$. Since $\mathcal{H}_{z}(\mathbf{x};\mathbf{w}_{z})$ has locally Lipschitz gradient, there exists $K_{\epsilon}^{1}$ such that if $\|\mathbf{w}_{1z}\|_{2},\|\mathbf{w}_{2z}\|_{2}\leq\tilde{B}_{\epsilon}$, then
$\left\|\mathcal{J}_{z}(\mathbf{X};\mathbf{w}_{1z})-\mathcal{J}_{z}(\mathbf{X};\mathbf{w}_{2z})\right\|_{2}\leq K^{1}_{\epsilon}\|\mathbf{w}_{1z}-\mathbf{w}_{2z}\|_{2}.$ (63)
Similarly, by local Lipschitz property of $\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})$ there exists $K_{\epsilon}^{2}$ such that if $\|\mathbf{w}_{1n}\|_{2}\leq\tilde{W}_{\epsilon}$, then
@@
$\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{1n})-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})\right\|_{2}\leq K^{2}_{\epsilon}\|\mathbf{w}_{1n}-\overline{\mathbf{w}}_{n}\|_{2}.$ (64)
We further define
$\bar{\delta}^{2}=\min\left(1,\epsilon^{2},\frac{\epsilon^{2}}{2C_{\epsilon}},\frac{\epsilon}{4\alpha\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}+\beta\tilde{B}_{\epsilon}^{L})T_{\epsilon}e^{2T_{\epsilon}K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}}}\right).$ (65)
Now, for any $\delta\in(0,\overline{\delta})$, let $\{\mathbf{w}_{n}(t),\mathbf{w}_{z}(t)\}$ satisfy for all $t\geq 0$
$\begin{bmatrix}\dot{\mathbf{w}}_{n}\\ \dot{\mathbf{w}}_{z}\end{bmatrix}=\begin{bmatrix}\mathcal{J}_{n}(\mathbf{X};\mathbf{w}_{n})^{\top}(\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z}))\\ \mathcal{J}_{z}(\mathbf{X};\mathbf{w}_{z})^{\top}(\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z}))\end{bmatrix},\begin{bmatrix}{\mathbf{w}}_{n}(0)\\ {\mathbf{w}_{z}}(0)\end{bmatrix}=\begin{bmatrix}\overline{\mathbf{w}}_{n}+\delta\mathbf{p}_{n}\\ \overline{\mathbf{w}}_{z}+\delta\mathbf{p}_{z}\end{bmatrix},$ (74)
where $\mathbf{p}_{n}$ and $\mathbf{p}_{z}$ are unit norm vectors. We define
$\mathbf{s}_{n}(t)=\mathbf{w}_{n}\left(\frac{t}{\delta^{L-2}}\right)\text{ and }\mathbf{s}_{z}(t)=\frac{1}{\delta}\mathbf{w}_{z}\left(\frac{t}{\delta^{L-2}}\right),$
where note that in contrast to $\mathbf{s}_{z}(t)$, $\mathbf{s}_{n}(t)$ does not have a factor of $1/\delta$. Then,
$\frac{d\mathbf{s}_{n}}{dt}$ $=\frac{1}{\delta^{L-2}}\dot{\mathbf{w}}_{n}\left(\frac{t}{\delta^{L-2}}\right)$
$=\frac{1}{\delta^{L-2}}\mathcal{J}_{n}\left(\mathbf{X};\mathbf{w}_{n}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{w}_{n}\left(\frac{t}{\delta^{L-2}}\right)\right)-\mathcal{H}_{z}\left(\mathbf{X};\mathbf{w}_{z}\left(\frac{t}{\delta^{L-2}}\right)\right)\right)$
$=\frac{1}{\delta^{L-2}}\mathcal{J}_{n}\left(\mathbf{X};\mathbf{s}_{n}\left(t\right)\right)^{\top}(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{s}_{n}\left(t\right)\right)-\delta^{L}\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\left(t\right)\right)),$ (75)
and
$\frac{d\mathbf{s}_{z}}{dt}$ $=\frac{1}{\delta^{L-1}}\dot{\mathbf{w}}_{z}\left(\frac{t}{\delta^{L-2}}\right)$
$=\frac{1}{\delta^{L-1}}\mathcal{J}_{z}\left(\mathbf{X};\mathbf{w}_{z}\left(\frac{t}{\delta^{L-2}}\right)\right)^{\top}\left(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{w}_{n}\left(\frac{t}{\delta^{L-2}}\right)\right)-\mathcal{H}_{z}\left(\mathbf{X};\mathbf{w}_{z}\left(\frac{t}{\delta^{L-2}}\right)\right)\right)$
$=\mathcal{J}_{z}\left(\mathbf{X};\mathbf{s}_{z}\left(t\right)\right)^{\top}(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{s}_{n}\left(t\right)\right)-\delta^{L}\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\left(t\right)\right)),$ (76)
where in the above set of equalities we mainly used $(L-1)$-homogeneity of the Jacobians and $L$-homogeneity of the neural network. We next define
$\mathbf{Z}(t)=\max(\|\mathbf{s}_{n}(t)-\overline{\mathbf{w}}_{n}\|_{2}^{2},\|\mathbf{s}_{z}(t)-\mathbf{u}(t)\|_{2}^{2}).$
Next, we will show that $\mathbf{Z}(t)\leq\epsilon$, for all $t\in[0,T_{\epsilon}]$. Note that $\mathbf{Z}(0)=\delta^{2}<\epsilon^{2}\leq\gamma^{2}$. Define
$\overline{T}_{\delta}=\inf_{t\in[0,T_{\epsilon}]}\{t:\mathbf{Z}(t)=\epsilon^{2}\}.$
Here, $\overline{T}_{\delta}$ indicates the time when $\mathbf{Z}(t)$ becomes $\epsilon^{2}$ for the first time. Thus, for all $t\in[0,\overline{T}_{\delta}],$ $\mathbf{Z}(t)\leq\epsilon^{2}\leq\gamma^{2}$. Note that if we can show $\overline{T}_{\delta}=T_{\epsilon}$, then it would imply $\mathbf{Z}(t)\leq\epsilon$, for all $t\in[0,T_{\epsilon}]$. Since $\overline{T}_{\delta}\leq T_{\epsilon}$, we have that $\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon}$ and $\|\mathbf{s}_{z}(t)\|_{2}\leq B_{\epsilon}+\epsilon=\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}]$. Also, $\|\mathbf{s}_{n}(t)\|_{2}\leq\|\overline{\mathbf{w}}_{n}\|_{2}+\epsilon=\tilde{W}_{\epsilon}$. From (63) and (64), for all $t\in[0,\overline{T}_{\delta}]$, we have
$\left\|\mathcal{J}_{z}(\mathbf{X};\mathbf{u}(t))-\mathcal{J}_{z}(\mathbf{X};\mathbf{s}_{z}(t))\right\|_{2}\leq K^{1}_{\epsilon}\|\mathbf{u}(t)-\mathbf{s}_{z}(t)\|_{2},$ (77)
and

### v2

@@ C.1 Key Lemmata
We begin by noting that, for $l\in[L]$,
$\partial_{{\mathbf{W}}_{l}}\left(\sum_{i=1}^{n}y_{i}\mathcal{H}({\mathbf{x}}_{i};{\mathbf{W}}_{1},\cdots{\mathbf{W}}_{L})\right)\subseteq\sum_{i=1}^{n}\partial_{{\mathbf{W}}_{l}}\left(y_{i}\mathcal{H}({\mathbf{x}}_{i};{\mathbf{W}}_{1},\cdots{\mathbf{W}}_{L})\right).$
Thus, we need to prove
@@
$\partial_{{\mathbf{W}}_{l}}\left(y_{i}\mathcal{H}({\mathbf{x}}_{i};{\mathbf{W}}_{1},\cdots{\mathbf{W}}_{L})\right)\subseteq{\mathbf{e}}_{i}^{l}(\phi_{i}^{l-1})^{\top},\text{ for all }l\in[L].$
For $l=L$, the above equation is true since
$\partial_{{\mathbf{W}}_{L}}\left(y_{i}\mathcal{H}({\mathbf{x}}_{i};{\mathbf{W}}_{1},\cdots{\mathbf{W}}_{L})\right)=\partial_{{\mathbf{W}}_{L}}\left(y_{i}{\mathbf{W}}_{L}\phi_{i}^{L-1}\right)=y_{i}\left(\phi_{i}^{L-1}\right)^{\top}.$
For $l\in[L-1]$, we have
@@
$\partial_{{\mathbf{W}}_{l}}\left(y_{i}\mathcal{H}({\mathbf{x}}_{i};{\mathbf{W}}_{1},\cdots{\mathbf{W}}_{L})\right)=\partial_{{\mathbf{W}}_{l}}\left(y_{i}{\mathbf{W}}_{L}\sigma({\mathbf{W}}_{L-1}\cdots\sigma({\mathbf{W}}_{l}\phi_{i}^{l-1})\cdots)\right)\subseteq{\mathbf{e}}_{i}^{l}(\phi_{i}^{l-1})^{\top},$
where the last inclusion follows from repeatedly applying the chain rule. This completes the proof. ∎
We next restate and prove Lemma 5.
Lemma 17.
Let ${\mathcal{H}}$ be an $L$-layer feed-forward neural network whose output is as defined in eq. 9, where $L\geq 2$. Suppose $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})$ is a non-zero KKT points of the constrained NCF in eq. 10. If $\sigma(x)=\max(\alpha x,x)^{p}$, for some $p\in{\mathbb{N}}$ and $\alpha\in{\mathbb{R}}$, then
$\text{diag}\left(\overline{{\mathbf{W}}}_{l}\overline{{\mathbf{W}}}_{l}^{\top}\right)=p\cdot\text{diag}\left(\overline{{\mathbf{W}}}_{l+1}^{\top}\overline{{\mathbf{W}}}_{l+1}\right),\text{ for all }l\in[L-1],$ (62)
and, if $\sigma(x)=x$, then
$\overline{{\mathbf{W}}}_{l}\overline{{\mathbf{W}}}_{l}^{\top}=\overline{{\mathbf{W}}}_{l+1}^{\top}\overline{{\mathbf{W}}}_{l+1},\text{ for all }l\in[L-1].$ (63)
Furthermore, if $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L-1},\overline{{\mathbf{W}}}_{L})=({\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top},\cdots,{\mathbf{a}}_{L-1}{\mathbf{b}}_{L-1}^{\top},\overline{{\mathbf{w}}}^{\top}),$ where $\|{\mathbf{a}}_{l}\|_{2}=\|{\mathbf{b}}_{l}\|_{2}$, for all $l\in[L-1]$, then $|{\mathbf{a}}_{l}|=p^{1/4}|{\mathbf{b}}_{l+1}|$, for all $l\in[L-2]$, $|{\mathbf{a}}_{L-1}|=(p\hat{p})^{1/4}|\overline{{\mathbf{w}}}|$, $\|{\mathbf{a}}_{l}\|_{2}^{4}=p^{L-l}/\hat{p},$ for all $l\in[L-1],$ and $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/\hat{p}$, where $\hat{p}=p^{L-1}+p^{L-2}+\cdots+1.$
Proof.
By KKT condition, for $l\in[L-1]$, we have
@@
$\mathbf{0}\in\partial_{{\mathbf{W}}_{l}}\left(\sum_{i=1}^{n}y_{i}\mathcal{H}({\mathbf{x}}_{i};\overline{{\mathbf{W}}}_{1},\cdots\overline{{\mathbf{W}}}_{L})\right)+\lambda\overline{{\mathbf{W}}}_{l}$ (64)
Multiplying the above equation by $\overline{{\mathbf{W}}}_{l}^{\top}$ from the right, and using Lemma 16, we have
$\mathbf{0}\in\sum_{i=1}^{n}{\mathbf{e}}_{i}^{l}(\phi_{i}^{l-1})^{\top}\overline{{\mathbf{W}}}_{l}^{\top}+\lambda\overline{{\mathbf{W}}}_{l}\overline{{\mathbf{W}}}_{l}^{\top}$ (65)
Since $\overline{{\mathbf{W}}}_{l}\phi_{i}^{l-1}={\mathbf{h}}_{i}^{l}$, we have
$\mathbf{0}\in\sum_{i=1}^{n}{\mathbf{e}}_{i}^{l}({\mathbf{h}}_{i}^{l})^{\top}+\lambda\overline{{\mathbf{W}}}_{l}\overline{{\mathbf{W}}}_{l}^{\top}.$ (66)
Since ${\mathbf{e}}_{i}^{l}={\mathbf{A}}_{i}^{l}\overline{{\mathbf{W}}}_{l+1}^{\top}\cdots{\mathbf{A}}_{i}^{L-1}\overline{{\mathbf{W}}}_{L}^{\top}y_{i}$, if we let ${\mathbf{q}}_{i}=\overline{{\mathbf{W}}}_{l+1}^{\top}\cdots{\mathbf{A}}_{i}^{L-1}\overline{{\mathbf{W}}}_{L}^{\top}y_{i}$, then we have
$\mathbf{0}\in\sum_{i=1}^{n}{\mathbf{A}}_{i}^{l}{\mathbf{q}}_{i}({\mathbf{h}}_{i}^{l})^{\top}+\lambda\overline{{\mathbf{W}}}_{l}\overline{{\mathbf{W}}}_{l}^{\top}.$ (67)
Next, similarly using the KKT condition for ${\mathbf{W}}_{l+1}$, we have
$\mathbf{0}\in\overline{{\mathbf{W}}}_{l+1}^{\top}\sum_{i=1}^{n}{\mathbf{e}}_{i}^{l+1}(\phi_{i}^{l})^{\top}+\lambda\overline{{\mathbf{W}}}_{l+1}^{\top}\overline{{\mathbf{W}}}_{l+1}.$ (68)
Expanding ${\mathbf{e}}_{i}^{l+1}$, we have
$\mathbf{0}\in\sum_{i=1}^{n}\overline{{\mathbf{W}}}_{l+1}^{\top}{\mathbf{A}}_{i}^{l+1}\overline{{\mathbf{W}}}_{l+2}^{\top}\cdots{\mathbf{A}}_{i}^{L-1}\overline{{\mathbf{W}}}_{L}^{\top}{\mathbf{y}}_{i}^{\top}(\phi_{i}^{l})^{\top}+\lambda\overline{{\mathbf{W}}}_{l+1}^{\top}\overline{{\mathbf{W}}}_{l+1},$ (69)
which implies
$\mathbf{0}\in\sum_{i=1}^{n}{\mathbf{q}}_{i}(\phi_{i}^{l})^{\top}+\lambda\overline{{\mathbf{W}}}_{l+1}^{\top}\overline{{\mathbf{W}}}_{l+1}.$ (70)
We next show that
$\text{diag}({\mathbf{A}}_{i}^{l}{\mathbf{q}}_{i}({\mathbf{h}}_{i}^{l})^{\top})=p\text{diag}({\mathbf{q}}_{i}(\phi_{i}^{l})^{\top}),\forall i\in[n]\text{ and }l\in[L-1],$
which will prove the first part of the theorem. To show this, first note that since ${\mathbf{A}}_{i}^{l}$ is a diagonal matrix, and ${\mathbf{q}}_{i},{\mathbf{h}}_{i}^{l}$ are vectors, we have
$\text{diag}({\mathbf{A}}_{i}^{l}{\mathbf{q}}_{i}({\mathbf{h}}_{i}^{l})^{\top})={\mathbf{A}}_{i}^{l}\text{diag}({\mathbf{q}}_{i})\text{diag}({\mathbf{h}}_{i}^{l})$ $=\text{diag}({\mathbf{q}}_{i}){\mathbf{A}}_{i}^{l}\text{diag}({\mathbf{h}}_{i}^{l})$
$=\text{diag}({\mathbf{q}}_{i})\text{diag}(\sigma^{\prime}({\mathbf{h}}_{i}^{l}))\text{diag}({\mathbf{h}}_{i}^{l})$
$=p\text{diag}({\mathbf{q}}_{i})\text{diag}(\sigma({\mathbf{h}}_{i}^{l}))$
$=p\text{diag}({\mathbf{q}}_{i})\text{diag}(\phi_{i}^{l})=p\text{diag}({\mathbf{q}}_{i}(\phi_{i}^{l})^{\top}),$
where in the third equality, we used the definition of ${\mathbf{A}}_{i}^{l}$ and the fourth equality follows from the fact $\sigma^{\prime}(x)x=p\sigma(x)$.
For $\sigma(x)=x$, note that ${\mathbf{A}}_{i}^{l}$ will be identity and $\phi_{i}^{l}={\mathbf{h}}_{i}^{l}$, for all $i\in[n]$ and $l\in[L-1]$. Hence,
${\mathbf{A}}_{i}^{l}{\mathbf{q}}_{i}({\mathbf{h}}_{i}^{l})^{\top}={\mathbf{q}}_{i}(\phi_{i}^{l})^{\top},\forall i\in[n]\text{ and }l\in[L-1],$
which proves eq. 63.
We next prove the second part. From eq. 62, we have
$\|\overline{{\mathbf{W}}}_{1}\|_{F}^{2}=p\|\overline{{\mathbf{W}}}_{2}\|_{F}^{2}=\cdots=p^{L-1}\|\overline{{\mathbf{W}}}_{L}\|_{F}^{2}.$
Since
$\|\overline{{\mathbf{W}}}_{1}\|_{F}^{2}+\cdots\|\overline{{\mathbf{W}}}_{L}\|_{F}^{2}=1,$
if we define $\hat{p}=p^{L-1}+p^{L-2}+\cdots+1$, then,
$1=\|\overline{{\mathbf{W}}}_{L}\|_{F}^{2}(p^{L-1}+p^{L-2}+\cdots+1),$
which implies $\|\overline{{\mathbf{W}}}_{l}\|_{F}^{2}=p^{L-l}/\hat{p}$, for all $l\in[L]$. Next, since $\|{\mathbf{a}}_{l}\|_{2}=\|{\mathbf{b}}_{l}\|_{2}$, we further get $p^{L-l}/\hat{p}=\|{\mathbf{a}}_{l}\|_{2}^{4},$ for all $l\in[L-1]$ and $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/\hat{p}.$
Now, using eq. 62, for all $l\in[L-2]$, we have
$\text{diag}\left(\|{\mathbf{b}}_{l}\|_{2}^{2}{\mathbf{a}}_{l}{\mathbf{a}}_{l}^{\top}\right)=p\cdot\text{diag}\left(\|{\mathbf{a}}_{l+1}\|_{2}^{2}{\mathbf{b}}_{l+1}{\mathbf{b}}_{l+1}^{\top}\right),\text{ and }\text{diag}\left(\|{\mathbf{b}}_{L-1}\|_{2}^{2}{\mathbf{a}}_{L-1}{\mathbf{a}}_{L-1}^{\top}\right)=p\cdot\text{diag}\left(\overline{{\mathbf{w}}}\overline{{\mathbf{w}}}^{\top}\right)$
which implies
$|{\mathbf{a}}_{l}|={p}^{1/4}|{\mathbf{b}}_{l+1}|\text{ and }|{\mathbf{a}}_{L-1}|={(p\hat{p})}^{1/4}|\overline{{\mathbf{w}}}|,$ (71)
where the last equality uses $p^{L-l}/\hat{p}=\|{\mathbf{b}}_{L-1}\|_{2}^{4}.$ Hence, the proof is complete. ∎
C.2 Proof of Theorem 3
From Lemma 5, we know $\|\overline{{\mathbf{w}}}\|_{2}^{2}=1/L$ and
$\|{\mathbf{a}}_{l}\|_{2}^{2}=\|{\mathbf{b}}_{l}\|_{2}^{2}=\frac{1}{\sqrt{L}},\text{ for all }l\in[L-1].$ (72)
For the remaining proof, we handle each of the cases individually, and begin by $\alpha=1$.
Case 1 $(\alpha=1):$ We begin by proving the “only if” part. From Lemma 5, we have
$\overline{{\mathbf{W}}}_{l}\overline{{\mathbf{W}}}_{l}^{\top}=\overline{{\mathbf{W}}}_{l+1}^{\top}\overline{{\mathbf{W}}}_{l+1},\forall l\in[L-1],$ (73)
which implies, for $l\in[L-2]$,
$\|{\mathbf{b}}_{l}\|_{2}^{2}{\mathbf{a}}_{l}{\mathbf{a}}_{l}^{\top}=\|{\mathbf{a}}_{l+1}\|_{2}^{2}{\mathbf{b}}_{l+1}{\mathbf{b}}_{l+1}^{\top},\text{ and }\|{\mathbf{b}}_{L-1}\|_{2}^{2}{\mathbf{a}}_{L-1}{\mathbf{a}}_{L-1}^{\top}=\overline{{\mathbf{w}}}\overline{{\mathbf{w}}}^{\top}.$
Combining the above equality with eq. 72 we get
$q_{l}{\mathbf{a}}_{l}={\mathbf{b}}_{l+1},\forall l\in[L-2],\text{ and }q_{L-1}{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}},$
where $q_{l}\in\{-1,1\}$. Hence,
$\overline{{\mathbf{W}}}_{1}={\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}\text{ and }\overline{{\mathbf{W}}}_{l}=q_{l-1}{\mathbf{a}}_{l}{\mathbf{a}}_{l-1}^{\top},\text{ if }2\leq l\leq L-1.$
Now, since $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})$ is a non-zero KKT point of eq. 10, using Lemma 11, we have